Calculate 3(22 - 6) – 5(4 ÷ 2) – (-15 + 3).
Found 2 solutions by rfer, jenni.fowler021510:
Answer by rfer(16322)

3*16-5*2+12
48-10+12=50

Answer by jenni.fowler021510(3)

3(18)-5(2)+15-3=
54 -10 +12=
44 +12=
56
so the answer is 56